OpenStudy (anonymous):

How do you factor ones with the power of 6?

OpenStudy (anonymous):

\[x ^{6}-y ^{6}\]

OpenStudy (lgbasallote):

difference of two squares or difference of two cubes ;)

mathslover (mathslover):

\[\Huge{x^6-y^6}\] \[\Huge{(x^3)^2-(y^3)^2}\] \[\Huge{(x^3+y^3)(x^3-y^3)}\] \[\large{[(x+y)(x^2-xy+y^2)][(x-y)(x^2+xy+y^2)]}\] \[\large{(X^2-y^2)(x^2-xy+y^2)(x^2+xy+y^2)}\]

OpenStudy (lgbasallote):

\[(x^2)^3 - (y^2)^3\] or \[(x^3)^2 - (y^3)^2\]

mathslover (mathslover):

\[\Huge{x^2-y^2=(x+y)(x-y)}\] \[\Huge{x^3-y^3=(x-y)(x^2+y^2+xy)}\] \[\Huge{x^3+y^3=(x+y)(x^2+y^2-xy)}\]

mathslover (mathslover):

these are the three identites that we used here